OpenStudy (anonymous):

What are the amplitude, period, and midline of f(x) = 5 sin(x − pi) + 3? PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Amplitude is the number in front of sin so your amplitude would be 5. To find the period you use the formula \[\frac{ 2\Pi }{ b }\]

OpenStudy (sidsiddhartha):

first lets consider a general function \[f(x) = A \sin(Bx - C) + D \] fot this function AMPLITUDE =A The period of the function is: 2π/B The phase shift of the function is: C/B so now just compare ur function with this one

OpenStudy (anonymous):

ok so 5, 2pi/1, and pi/x???

OpenStudy (anonymous):

Amplitude: 5; period: 𝜋; midline: y = 3 Amplitude: 3; period: pi over 2; midline: y = 5 Amplitude: 3; period: pi over 4; midline: y = 5 Amplitude: 5; period: 𝜋; midline: y = 3 those are the answer choices...

OpenStudy (anonymous):

idk why a and d are the same...

OpenStudy (sidsiddhartha):

yeah ine one of them period should be =2*(pi)

OpenStudy (sidsiddhartha):

Amplitude: 5; period: 2𝜋; midline: y = 3 this should be the answer

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

yah but wich one is the answer because a and d are the same...

OpenStudy (anonymous):

because i got that too but a and d are the same.....

OpenStudy (anonymous):

Amplitude: 5; period: 𝜋; midline: y = 3 Amplitude: 3; period: pi over 2; midline: y = 5 Amplitude: 3; period: pi over 4; midline: y = 5 Amplitude: 5; period: 𝜋; midline: y = 3 AAnswers right there

OpenStudy (sidsiddhartha):

yeah that make no option correct lol i think there should be some misprint

OpenStudy (anonymous):

yah.. ik what to do...

OpenStudy (anonymous):

idk*

OpenStudy (sidsiddhartha):

look at this the period is definately 2pi http://www.wolframalpha.com/input/?i=period+f%28x%29+%3D+5+sin%28x+%E2%88%92+pi%29+%2B+3

OpenStudy (anonymous):

yah... but i need the midline.. .what would the midline be?

OpenStudy (sidsiddhartha):

midline is the value of D in the above equation and it is 3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i just put a.. can you help with more... i have like 15 more... ill give medals

OpenStudy (sidsiddhartha):

yeah will try :)